A purely competitive wheat farmer can sell any wheat he grows for $30 per bushel. His five acres of land show diminishing returns, because some are better suited for wheat production than others. The first acre can produce 1000 bushels of wheat, the second acre 900, the third 800, and so on. In the table given below answer the following questions. How many bushels will each of the farmer’s five acres produce? How much revenue will each acre generate? What are the TR and MR for each acre?
1
Expert's answer
2013-07-22T09:15:32-0400
Apurely competitive wheat farmer can sell any wheat he grows for $30 per bushel.
His five acres of land show diminishing returns, because some are better suited
for wheat production than others. The first acre can produce1000 bushels of
wheat, the second acre 900, the third 800, and so on.
1) How many bushels will each of the farmer’s five acres produce?
Acre Quantity
Total quantity Revenue per acre
TR MR
1 acre 1000 bushels 1000

30,000 30,000
-
2 acre 900 bushels 1900

27,000 57,000
30
3 acre 800 bushels 2700

24,000 81,000
30
4 acre 700 bushels 3400

21,000 102,000
30
5 acre 600 bushels 4000

18,000 120,000
30
2) How much revenue will each acre generate?
Revenue per acre = P*Q, P = $30
3) What are the TR and MR for each acre?
TR = Q*Revenue per acre
MR = (TR2 - TR1)/(Q2 - Q1) = P = $30